OpenStudy (anonymous):

mimi is four years older than ronald. five years ago she was twice as old as he was. find their ages now

OpenStudy (henryblah):

Let mimi's age now be x. Ronald now to be y. y+4=x x-5=2(y-5) Solve simultaneously

OpenStudy (radar):

Let R = Ronald's age now. then, Mimi is R+4 now now for the 5 years ago scenario: R+4 -5 = 2(R-5) R-1=2R-10 -R=-9 R=9 Ronald age now 9+4=13 Mimi age now 5 yrs ago they were 4 and 8 respectively, note that Mimi is twice as old as Ronald. Same answer as Henryblah's method but only using a single variable.

OpenStudy (anonymous):

Solve the following equations for m and r:\[\left\{m-4=r,\frac{m-5}{r-5}=2\right\}\]\[\{m=13,r=9\} \]

OpenStudy (anonymous):

@radar I gave you a medal because you were there first with the solutions.

OpenStudy (henryblah):

Hmm, I don't like giving answers out. I just give helpful prods :)

OpenStudy (radar):

@Henryblah, Sometimes a prob needs a push lol